# High-dimensional limit theorems for SGD: Momentum and Adaptive Step-sizes

### v1

@@ 1. Introduction
The fixed-dimensional asymptotic theory of SGD is classical [robbins1951stochastic, mcleish1976functional, kushner1984asymptotic]. In this regime, the small step-size limit of online SGD converges to the gradient flow on the population loss. More recently, in the high dimensional regime, there have been numerous works on scaling limits where the dimension tends to infinity for specific problems, including linear regression [wang2017scaling, paquette2021sgd, paquette2024homogenization], online PCA [li2016online, wang2017scaling], single-index models [saad1995dynamics, saad1995line, goldt2019dynamics, veiga2022phase], multi-index models [collins2024hitting], as well as a unifying framework that further extends to diffusive dynamics [ben2022high, benarous2024high]. A key insight from the high-dimensional perspective is that when the dimension scales up as the step-size tends to zero, there exists a critical scaling regime of the step-size in which high-dimensional effects yield distinct dynamics. The fixed-dimensional dynamics are recovered with sub-critical step-size, but at the critical scaling, there are additional corrections, such as the “population corrector,” in addition to the gradient flow drift.
For SGD with Polyak or Nesterov momentum [polyak1964some, nesterov269method, sutskever2013importance], there are various recent works in the fixed dimensional setting. Some develop similar continuous-time ballistic limits [su2016differential, kovachki2021continuous, wilson2021lyapunov, feng2023anytime] with a related line of work examining continuous-time limits under heavy-tailed noise [simsekli2020fractional, dang2025algorithmic]. There is also a well-established body of literature characterizing the benefits of SGD-M over online SGD, such as improved convergence rates, reduced escape times from saddle points, and implicit regularization effects [liu2021diffusion, wang2021escaping, cowsik2022flatter, gess2023exponential, ghosh2023implicit, dang2025algorithmic]. Conversely, the dynamics of several momentum-based stochastic algorithms were studied for linear regression [paquette2021dynamics].
Preconditioned SGD methods incorporating adaptive step-sizes [duchi2011adaptive, hinton2012rmsprop, zeiler2012adadelta, kingma2014adam, yu2017block, shazeer2018adafactor] are also of theoretical interest, particularly for explaining their empirical success in high-dimensional, non-convex optimization. In fixed-dimensional settings, notable works addressed similar questions regarding convergence behaviour and continuous-time dynamics [da2018general, reddi2019convergence, barakat2021convergence, ma2022qualitative, malladi2022sdes, ghosh2023implicit]. The preconditioner we consider in Section 3, gradient normalization, was first introduced to mitigate exploding gradients in deep sequence models [mikolov2012statistical, pascanu2013difficulty]. Since then, gradient normalization has remained a central motif in adaptive step-size design, supporting both layerwise and component-wise control [duchi2011adaptive, hinton2012rmsprop, kingma2014adam, yu2017block].
@@ 1.1. Contributions
We extend the effective dynamics framework [benarous2024high] to derive the limiting dynamics of SGD with Polyak momentum (SGD-M) and of SGD with adaptive step-sizes. For a fixed step-size in the critical scaling regime, the limiting dynamics of SGD-M accentuates the emergent “population corrector” observed in online SGD, potentially steering the dynamics further away from the population gradient. However, for any given step-size of SGD-M, there exists a corresponding step-size for online SGD that yields the same effective dynamics, up to a time rescaling, establishing an equivalence between the algorithms. Furthermore, in our examples, we compare the evolution of SGD-M with that of a variant of online SGD equipped with a simple adaptive step-size that reduces to unit norm gradients [mikolov2012statistical, pascanu2013difficulty]; we refer to this variant as “SGD-U” in subsequent sections.
While our results are broadly applicable, we illustrate them on two canonical problems in high dimensional inference: spiked tensor PCA [johnstone2001distribution, peche2006largest, montanari2014statistical] and single index models [bishop2006pattern, hastie2009elements, barbier2019optimal]. In both settings, we confirm that in the critical scaling regime, SGD-M can at best replicate the behaviour of online SGD. By contrast, we demonstrate a substantial difference between online SGD and SGD-U: the latter exhibits dynamics that converge to superior fixed points (i.e. closer to the population minimum) and significantly broadens the range of admissible step-sizes that ensure convergence to these solutions. The admissibility of larger step-sizes that maintain stable dynamics and yield high-quality solutions provides a rigourous demonstration of how preconditioners can mitigate high-dimensional effects and empirical phenomena such as exploding or vanishing gradients. Thus, our framework offers theoretical justification for the empirical motivations underlying the design of early preconditioners.
@@ 2. Main Result
We consider the following online learning setup. Suppose we are given an i.i.d. data sequence $\boldsymbol{y}_{1},\boldsymbol{y}_{2},\ldots$ taking values in $\mathcal{Y}\subseteq\mathbb{R}^{d_{n}}$ with law $\mu\in\mathcal{M}_{1}(\mathbb{R}^{d_{n}})$ and loss $L:\mathcal{X}_{n}\times\mathcal{Y}_{n}\to\mathbb{R}$ where $\mathcal{X}_{n}\subseteq\mathbb{R}^{p_{n}}$ is the parameter space. Consider online SGD with learning rate $\delta_{n}>0$ and momentum rate $\beta\in[0,1)$, given by
(1) $\boldsymbol{p}_{\ell}$ $=\beta\boldsymbol{p}_{\ell-1}-\delta_{n}\,\nabla L(\boldsymbol{x}_{\ell-1},\boldsymbol{y}_{\ell})$
@@
To proceed, we state two assumptions: $\delta_{n}$-localizability and $\delta_{n}$-closability. The first, which we present next, provides an upper bound on the learning rate in terms of the regularity of the summary statistics and data distribution. These bounds ensure tightness of trajectories of the summary statistics.
@@
A quadruple $(\boldsymbol{u}_{n},\nabla L_{n},P_{n},\beta)$ is $\delta_{n}$-localizable with localizing sequence $(E_{K})_{K}$ if there is an exhaustion by compacts $(E_{K})_{K}$ of $\mathbb{R}^{k}$, and constants $C_{K}$ (independent of $n$) such that
@@
    $\max_{i}\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\|\nabla^{2}u_{i}^{n}\|_{\mathrm{op}}\leq C_{K}\cdot\delta_{n}^{-1/2}$, and $\max_{i}\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\|\nabla^{3}u_{i}^{n}\|_{\mathrm{op}}\leq C_{K}$;
@@
    $\max_{i}\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\mathbb{E}[\langle\nabla H,\nabla u_{i}^{n}\rangle^{4}]\leq C_{K}\delta_{n}^{-2}$, and
    $\max_{i}\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\mathbb{E}[\langle\nabla^{2}u_{i}^{n}(x),\nabla H(x,y)\otimes\nabla H(x,y)-V(x)\rangle^{2}]=o(\delta_{n}^{-3})$, and
    $\max_{i}\sup_{x_{1},x_{2}\in\boldsymbol{u}_{n}^{-1}(E_{K})}\mathbb{E}_{y_{1}\perp y_{2}}[\langle\nabla^{2}u_{i}^{n}(x_{1}),\nabla H(x_{1},y_{1})\otimes\nabla H(x_{2},y_{2})\rangle^{2}]=o(\delta_{n}^{-3})$
@@
Compared to the same definitions in the online SGD case [benarous2024high], we now require an additional item in (3) in Definition 2.1 that controls the correlation between the random fluctuations $\nabla H(x_{1},Y_{i})$, $\nabla H(x_{2},Y_{j})$ when $i\neq j$. This item is of a similar flavor as the preceding item in (3). Additionally, for Definition 2.2, we note that the two operators scale differently in $\beta$. We revisit this point in Remark 2.4 after presenting our main result below.
@@
Taking $\beta=0$ recovers the original result for online SGD [benarous2024high]. The proof is presented in Section 4.
@@
As a consequence of Remark 2.4, any trajectory produced by SGD-M over some time interval $[0,T]$ can be replicated by online SGD after a time change where both the step-size and time horizon are rescaled by $(1-\beta)^{-1}$. The effective number of iterations $T/\delta$ remains unchanged, confirming a similar observation, for example, in the fixed-dimensional deterministic regime [kovachki2021continuous], and the case of high-dimensional linear regression [paquette2021dynamics].
@@ 2.1. Scaling Limits of Adaptive step-sizes
which serve as the analogue to $\nabla\Phi$ and $\nabla H$ as defined before. Provided $\nabla\tilde{H}$ and $\nabla\tilde{\Phi}$ satisfy Definition 2.1 and 2.2, then Theorem 2.3 with $\beta=0$ extends naturally to the preconditioned case. Hence, we can rigorously derive scaling limits for this family of pre-conditioners under this same framework, which we demonstrate next.
@@ 3.1. Spiked Matrix and Tensor PCA
Suppose we are given i.i.d. samples of data of the form
$Y^{\ell}=\lambda v^{\otimes k}+W^{\ell}$
where $W^{\ell}$ are i.i.d. $k$-tensors with Gaussian entries, $v\in\mathbb{R}^{n}$ is a unit vector, and $\lambda=\lambda_{n}>0$ is the signal-to-noise ratio. Our goal is to infer the direction $v$ from the noisy sample.
@@
to optimize parameter vector $x$. We consider the following pair of summary statistics $\boldsymbol{u}_{n}=(m,r_{\perp}^{2})$ with
$m=m(x):=\langle x,v\rangle\quad\text{ and }\quad r_{\perp}^{2}=r_{\perp}^{2}(x):=\|x-mv\|^{2}=\|x\|^{2}-m^{2}$
For notational brevity we also define $R^{2}=m^{2}+r_{\perp}^{2}=\|x\|^{2}$ and set $r^{2}:=r^{2}_{\perp}$. For both cases we take the critical scaling of $\delta_{n}=c_{\delta}/n$ where $c_{\delta}>0$ is a constant. Since our goal is to estimate the direction $v$, we wish to optimize the ratio $|\frac{m}{R}|$, representing the magnitude of the cosine ratio between $x$ and $v$.
@@
(5) $\text{d}m=\frac{1}{1-\beta}2m\left(\lambda km^{k-2}-kR^{2(k-1)}\right)\text{d}t,\quad\text{d}r_{\perp}^{2}=-\frac{4kR^{2(k-1)}}{(1-\beta)^{2}}\left(r_{\perp}^{2}(1-\beta)-c_{\delta}\right)\text{d}t$
@@
We analyze the fixed points for both systems in the presented order. We find that the dynamics yield fixed points away from the axis $m=0$ (i.e. when $x$ and $v$ are not aligned) only if the signal $\lambda$ exceeds a critical value, say $\lambda_{\text{crit}}(k,\beta,c_{\delta})$, that depends on the tensor order, $k$, step-size, $c_{\delta}$ and $\beta$. For both SGD-M and SGD-U, we can show (see Section 5.2) that the critical $\lambda$ for the respective methods are
@@
In Figure 1, we provide simulations for Matrix PCA (i.e. $k=2$) with fixed $c_{\delta}=1$ to demonstrate how SGD-U and SGD-M with different values for $\beta$ and $\lambda$ behave in the ballistic phase. We see that as $\lambda^{U}_{\text{crit}}(2,c_{\delta})<\lambda^{M}_{\text{crit}}(2,\beta,c_{\delta})$ for $\beta\geq 0$, there are values of $\lambda$ for which SGD-U is supercritical when SGD-M and online SGD are subcritical. This means that SGD-U is able to converge towards one of the fixed points away from $m=0$ for values of $\lambda$ where the other methods fail to do so. A careful analysis reveals different settings of $c_{\delta}$ and $\lambda$ for which SGD-M may converge to desirable critical points instead. For example, fixing $c_{\delta}$, the fixed points suggest that SGD-U is preferred whenever $\lambda>(1-\beta)^{2}/8$. Similarly, the emergence of fixed points away from $m=0$ depends on the choice of $c_{\delta}$ relative to $\lambda$. Indeed, one can verify that the largest $c_{\delta}$ for which non-trivial fixed points emerge is larger under SGD-U than SGD-M whenever $\lambda>(1-\beta)^{2}/8$. This suggests that there is a problem-specific critical $\lambda$ of $\lambda=1/8$ that dictates which of online SGD or SGD-U is preferred.
@@ 3.1.1. Diffusive Limits at the Equator ($m=0$)
(8) $\text{d}r^{2}_{\perp}$ $=-2\left(r^{k+1}-\frac{c_{\delta}}{2}\right)\,\text{d}t$
We see that $r^{2}_{\perp}$ now solves an autonomous ODE which converges exponentially to $c_{\delta}/2]^{2/(k+1)}$. In particular, when $k=2$ and as $t$ tends to $\infty$, the equation for $\tilde{m}$ behaves like
$\text{d}\tilde{m}=\frac{2\tilde{m}}{c_{\delta}^{1/3}}\left(\lambda-\left[\frac{c_{\delta}}{2}\right]^{2/3}\right)\,\text{d}t+\left(\frac{2\tilde{m}^{2}}{c_{\delta}^{1/3}}+c_{\delta}\right)^{1/2}\text{d}B_{t}$
Interestingly, in contrast to the dynamics for online SGD [benarous2024high, Proposition 3.4], (7) is no longer a standard OU process as the diffusion coefficient explicitly depends on $\tilde{m}$. However it is still reminiscent of one since, for example, the drift is mean-reverting or mean-repellent depending on the choice of $c_{\delta}$ relative to $\lambda$. Figure 2 presents several simulations of the diffusive limits for $\tilde{u}_{1}$ under different setups of SGD-M and SGD-U.

### v2

@@ 1. Introduction
The fixed-dimensional asymptotic theory of SGD is classical, dating back to [24, 33, 45]. In this regime, the small step-size limit of online SGD converges to the gradient flow on the population loss. More recently, in the high dimensional regime, there have been numerous works on scaling limits where the dimension tends to infinity for specific problems, including linear regression [37, 38, 56], online PCA [25, 56], single-index models [17, 47, 46, 54], and multi-index models [8]. A unifying framework for proving such scaling limits was introduced by [4, 5], including an extension to study diffusive dynamics. A key insight from the high-dimensional perspective is that when the dimension scales up as the step-size tends to zero, there exists a critical scaling regime of the step-size in which high-dimensional effects yield distinct dynamics. The fixed-dimensional dynamics are recovered with sub-critical step-size, but at the critical scaling, there are additional corrections, such as the “population corrector,” in addition to the gradient flow drift.
For SGD with momentum [36, 42, 53], recent works in the fixed dimensional setting consider similar continuous-time ballistic limits [13, 23, 52, 58]. There is also a well-established body of literature characterizing the benefits of SGD-M over online SGD, such as improved convergence rates, reduced escape times from saddle points, and implicit regularization effects [9, 11, 15, 16, 28, 57]. Conversely, in the high-dimensional setting, [14, 39] study the limiting dynamics of several momentum-based stochastic algorithms in the setting where they scale the momentum parameter accordingly.
Preconditioned SGD methods incorporating adaptive step-sizes [12, 19, 21, 48, 59, 60] are also of theoretical interest, particularly for explaining their empirical success in high-dimensional, non-convex optimization [62]. In fixed-dimensional settings, notable works addressed similar questions regarding convergence behaviour and continuous-time dynamics [2, 10, 16, 29, 31, 44]. The preconditioner we consider in Section 3, gradient normalization, was first introduced to mitigate exploding gradients in deep sequence models [34, 40]. Since then, there have been several works looking at the effectiveness of “clipped” gradient control [7, 22, 26, 61] as well as high-dimensional scaling limits for random least squares regression [32].
@@ 1.1. Contributions
We extend the framework of [5] to derive the high-dimensional limiting dynamics of SGD with Polyak momentum (SGD-M) and of SGD with adaptive step-sizes. We find that the critical step-size scalings are the same between these dynamics. Furthermore, we find that in the critical step-size regime, SGD-M amplifies the high-dimensional effects as compared to SGD, potentially steering the dynamics even further from the population gradient. More precisely, for any sequence of step-sizes in the critical scaling regime, the effective dynamics of SGD-M accentuates the aforementioned emergent “population corrector” observed in online SGD. However, for any given step-size of SGD-M, there exists a corresponding step-size for online SGD that yields the same effective dynamics, up to a time rescaling, establishing an equivalence between the algorithms.
In contrast, we find that training can be improved by incorporating pre-conditioning. To this end, we make use of our framework to compare SGD-M with a version of online SGD equipped with a simple adaptive step-size that restricts the gradients to have unit norm [34, 40]; we refer to this variant as “SGD-U” in subsequent sections. We prove, in two popular classes of learning tasks, that the SGD-U can exhibit dynamics with not only superior fixed points (i.e. closer to the population minimum) but can significantly broaden the range of admissible step-sizes that ensure convergence to these solutions. The admissibility of larger step-sizes that maintain stable dynamics and yield high-quality solutions provides a rigourous demonstration of how preconditioners can mitigate high-dimensional effects and empirical phenomena such as exploding or vanishing gradients. Thus, our framework offers rigorous justification for the empirical motivations underlying the design of early preconditioners.
While our framework is broadly applicable, we illustrate our results on two problems in high dimensional inference: spiked tensor PCA [20, 35, 41] and single index models [3, 6, 18].
@@ 2. Main Result
We consider the following online learning setup. Suppose we are given an i.i.d. data sequence $\boldsymbol{y}_{1},\boldsymbol{y}_{2},\ldots$ taking values in $\mathcal{Y}\subseteq\mathbb{R}^{d_{n}}$ with law $P_{n}\in\mathcal{M}_{1}(\mathbb{R}^{d_{n}})$ and loss $L:\mathcal{X}_{n}\times\mathcal{Y}_{n}\to\mathbb{R}$ where $\mathcal{X}_{n}\subseteq\mathbb{R}^{p_{n}}$ is the parameter space. Consider online SGD with learning rate $\delta_{n}>0$ and momentum rate $\beta\in[0,1)$, given by
(1) $\boldsymbol{p}_{\ell}$ $=\beta\boldsymbol{p}_{\ell-1}-\delta_{n}\,\nabla L_{n}(\boldsymbol{x}_{\ell-1},\boldsymbol{y}_{\ell})$
@@
To proceed, we state two assumptions: $\delta_{n}$-localizability and asymptotic closability. The first, which we present next, provides an upper bound on the learning rate in terms of the regularity of the summary statistics and data distribution. These bounds ensure tightness of trajectories of the summary statistics.
@@
A tuple $(\boldsymbol{u}_{n},\nabla L_{n},P_{n},\beta)$ is $\delta_{n}$-localizable with localizing sequence $(E_{K})_{K}$ if there is an exhaustion by compacts $(E_{K})_{K}$ of $\mathbb{R}^{k}$, and constants $C_{K}$ such that for any $i\leq k$,
@@
    $\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\|\nabla^{2}u_{i}^{n}\|_{\mathrm{op}}\leq C_{K}\cdot\delta_{n}^{-1/2}$, and $\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\|\nabla^{3}u_{i}^{n}\|_{\mathrm{op}}\leq C_{K}$;
@@
    $\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\mathbb{E}[\langle\nabla H,\nabla u_{i}^{n}\rangle^{4}]\leq C_{K}\delta_{n}^{-2}$, and
    $\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\mathbb{E}[\langle\nabla^{2}u_{i}^{n}(x),\nabla H(x,y)\otimes\nabla H(x,y)-V(x)\rangle^{2}]=o(\delta_{n}^{-3})$, and
    $\sup_{(x_{i})_{1}^{2}\in\boldsymbol{u}_{n}^{-1}(E_{K})}\mathbb{E}_{y_{1}\perp y_{2}}[\langle\nabla^{2}u_{i}^{n}(x_{1}),\nabla H(x_{1},y_{1})\otimes\nabla H(x_{2},y_{2})\rangle^{2}]=o(\delta_{n}^{-3})$ for $\beta>0$
@@
We refer the reader to [5] for a detailed discussion regarding the interpretations of Definition 2.1. In particular, these scaling assumptions are more general than the regularity assumptions commonly imposed in the literature, e.g., $L$-smoothness and convexity of the loss. These assumptions can be checked in a broad class of models, such as those studied here or in [5] and related works. For example, one can verify that Definition 2.1 holds for common models such as binary classification and Gaussian XOR classification with two-layer neural networks with ReLU activations [5].
Let us briefly compare the notions of localizability and asymptotically closability to the related notions in [5] . Compared to online SGD, we now require an additional item in Definition 2.1-(3) to control the correlation between the random fluctuations $\nabla H(x_{1},Y_{i})$, $\nabla H(x_{2},Y_{j})$ when $i\neq j$. Additionally, for Definition 2.2, we note that the two operators scale differently in $\beta$. We revisit this point in Remark 2.4 after presenting our main result below.
@@
Taking $\beta=0$ recovers the original result due to [5] for online SGD. The proof is presented in Appendix 6.
@@
As a consequence of Remark 2.4, any trajectory produced by SGD-M over some time interval $[0,T]$ can be replicated by online SGD after a time change where both the step-size and time horizon are rescaled by $(1-\beta)^{-1}$. The effective number of iterations $T/\delta$ remains unchanged, confirming a similar observation to both [23], who considers the fixed-dimensional deterministic regime, and [39], who considers the high-dimensional regime for the case of linear regression.
@@ 2.1. Scaling Limits of Adaptive step-sizes
which serve as the analogue to $\nabla\Phi$ and $\nabla H$ as defined before. Provided $\nabla\tilde{H}$ and $\nabla\tilde{\Phi}$ satisfy Definition 2.1 and 2.2, then Theorem 2.3 with $\beta=0$ extends naturally to the preconditioned case (see Appendix 5 for an extended discussion). Hence, we can rigorously derive scaling limits for this family of pre-conditioners under this same framework, which we demonstrate next.
@@ 3.1. Spiked Matrix and Tensor PCA
Suppose we are given i.i.d. samples of data of the form $Y^{\ell}=\lambda v^{\otimes k}+W^{\ell}$ where $W^{\ell}$ are i.i.d. $k$-tensors with Gaussian entries, $v\in\mathbb{R}^{n}$ is a unit vector, and $\lambda=\lambda_{n}>0$ is the signal-to-noise ratio. Our goal is to infer the direction $v$ from the noisy sample.
@@
to optimize parameter vector $x$. We consider the pair of summary statistics $\boldsymbol{u}_{n}=(m,r^{2})$ with
$m=m(x):=\langle x,v\rangle\quad\text{ and }\quad r^{2}=r^{2}(x):=\|x-mv\|^{2}=\|x\|^{2}-m^{2}$
We also define $R^{2}=m^{2}+r^{2}=\|x\|^{2}$. For both cases we take the critical scaling of $\delta_{n}=c_{\delta}/n$ where $c_{\delta}>0$ is a constant. Since our goal is to estimate the direction $v$, we wish to optimize the ratio $|\frac{m}{R}|$, representing the magnitude of the cosine ratio between $x$ and $v$.
@@
(5) $\text{d}m=\frac{1}{1-\beta}2m\left(\lambda km^{k-2}-kR^{2(k-1)}\right)\text{d}t,\quad\text{d}r^{2}=-\frac{4kR^{2(k-1)}}{(1-\beta)^{2}}\left(r^{2}(1-\beta)-c_{\delta}\right)\text{d}t$
@@
We verify localizability and closability for both SGD-M and SGD-U in Sections 7.1.1, 7.1.2 as part of the proof of Proposition 3.1. We analyze the fixed points for both systems in the presented order. We find that the dynamics yield fixed points away from the axis $m=0$ only if $\lambda$ exceeds a critical value, say $\lambda_{\text{crit}}(k,\beta,c_{\delta})$, that depends on the tensor order, $k$, step-size, $c_{\delta}$ and $\beta$. For both SGD-M and SGD-U, we can show (see Appendix 7.2) that the critical $\lambda$ for the respective methods are
@@
In Figure 1, we provide simulations for Matrix PCA (i.e. $k=2$) with fixed $c_{\delta}=1$ to demonstrate how SGD-U and SGD-M with different values for $\beta$ and $\lambda$ behave in the ballistic phase. We see that as $\lambda^{U}_{\text{crit}}(2,c_{\delta})<\lambda^{M}_{\text{crit}}(2,\beta,c_{\delta})$ for $\beta\geq 0$, there are values of $\lambda$ for which SGD-U is supercritical when SGD-M and online SGD are subcritical. This means that SGD-U is able to converge towards one of the fixed points away from $m=0$ for values of $\lambda$ where the other methods fail to do so. A careful analysis reveals different settings of $c_{\delta}$ and $\lambda$ for which SGD-M may converge to desirable critical points instead. For example, fixing $c_{\delta}$, the fixed points suggest that SGD-U is preferred whenever $\lambda>(1-\beta)^{2}/8$. Similarly, the emergence of fixed points away from $m=0$ depends on the choice of $c_{\delta}$ relative to $\lambda$. Indeed, one can verify that the largest $c_{\delta}$ for which non-trivial fixed points emerge is larger under SGD-U than SGD-M whenever $\lambda>(1-\beta)^{2}/8$. This suggests that there is a problem-specific critical $\lambda=1/8$ dictating which of online SGD or SGD-U is preferred.
@@ 3.1.1. Diffusive Limits at the Equator ($m=0$)
(8) $\text{d}r^{2}$ $=-2\left(r^{k+1}-\frac{c_{\delta}}{2}\right)\,\text{d}t$
We see that $r^{2}$ now solves an autonomous ODE which converges exponentially to $c_{\delta}/2]^{2/(k+1)}$. In particular, when $k=2$ and as $t$ tends to $\infty$, the equation for $\tilde{m}$ behaves like
$\text{d}\tilde{m}=\frac{2^{1/3}\tilde{m}}{c_{\delta}^{1/3}}\left(\lambda-\left[\frac{c_{\delta}}{2}\right]^{2/3}\right)\,\text{d}t+\left(\frac{2^{2/3}\tilde{m}^{2}}{c_{\delta}^{1/3}}+c_{\delta}\right)^{1/2}\text{d}B_{t}$
Interestingly, in contrast to the dynamics for online SGD [5, Proposition 3.4], (7) is no longer a standard OU process as the diffusion coefficient explicitly depends on $\tilde{m}$. However it is still reminiscent of one since, for example, the drift is mean-reverting or mean-repellent depending on the choice of $c_{\delta}$ relative to $\lambda$. Figure 2 presents several simulations of the diffusive limits for $\tilde{u}_{1}$ under different setups of SGD-M and SGD-U.